OpenStudy (anonymous):

how to integrate dy/dx=(5-x)/(x)??

OpenStudy (anonymous):

\[\frac{dy}{dx}=\frac{5-x}{x}=\frac{5}{x}-1\] Integrate term by term.

OpenStudy (anonymous):

I have to find the general solution, is that the answer?

OpenStudy (anonymous):

The general solution would be \(y=\cdots\), so you actually have to integrate first.

OpenStudy (anonymous):

can you help me to do that?

OpenStudy (anonymous):

\[y=\int\frac{dy}{dx}~dx=\int\left(\frac{5}{x}-1\right)~dx\]

OpenStudy (anonymous):

Yes ihave that already

OpenStudy (anonymous):

Log rule for the first term, power rule for the second. Just apply what you've (presumably) learned.

OpenStudy (anonymous):

I am confused still